OpenStudy (anonymous):

the sides of a triangle are in the ratio 5 : 12 : 13. what is the length of each side of the triangle if the perimeter of the triangle is 15 inches?

OpenStudy (anonymous):

The ratio would add up to a perimeter of 30, so your perimeter is 15; each side will be in the same ratio of 1:2 5* 1/2 = 2.5 inches 12 * 1/2 = 6 inches 13 * 1/2 = 6.5 inches 2.5 + 6 + 6.5 = 15 inches. (This does check)

OpenStudy (campbell_st):

use the unitary method to find the value of 1 part in the ratio. start with adding the parts in the ratio.... 5 + 12 + 13= the divide the perimeter by the total number of parts I unit = 15/total parts lastly, multiply each value in the ratio by the value of 1 unit hope it helps